The Corata Appliance Manufacturing Corporation manufactures two vacuum cleaners, the Standard and the Super. The following information was gathered about the two products:

                                                                                           Standard               Super

    Budgeted sales in units                                                 2700                       600

    Budgeted selling price                                                   $600                  $1,700

    Budgeted contribution margin per unit                   $600                   $1080

    Actual sales in units                                                      3000                     1400

    Actual selling price                                                        $650                  $1,680

What is the total sales-mix variance in terms of the contribution margin? (Round intermediary calculations to two decimal places.)

A.

$295,680 favorable

B.

$1,034,880 favorable

C.

$369,600 favorable

D.

$665,280 favorable
Answer #1

Answer is option A

A.

$295,680 favorable

Actual units of all products sold

Actual sales-mix % - Budgeted sales-mix %

Budgeted CM per unit

Sales-mix variance

Standard

4400

X

(0.68-0.82)

X

600

=

369600 U

super

4400

X

(0.32-0.18)

X

1080

=

665280 F

$295680 F

Actual sales mix

Standard = 3000/4400 = 0.68

Super = 1400/4400 = 0.32

Budgeted sales mix

Standard = 2700/3300 = 0.82

Super = 600/3300 = 0.18